THE STRUCTURE OF ETHYNYLFERROCENE USING MICROWAVE SPECTROSCOPY^{a}

RANGA SUBRAMANIAN, CHANDANA KARUNATILAKA, KRISTEN KECK and \underline{STEPHEN KUKOLICH}, Department of Chemistry, University of Arizona, Tucson, AZ 85721.

Structural parameters for ETHYNYLFERROCENE and other singlysubstituted ferrocenes have been determined, using microwave spectroscopy and DFT calculations. For ethynylferrocene twenty-four rotational constants have been determined by fitting the measured frequencies of a- and b- type transitions, for various isotopomers. The distance between the Fe-atom and the center of the cyclopentadienyl ring is 1.648(5) Åand the distance between the carbon atoms of the cyclopentadienyl ring is r_{CC} = 1.432(2)Å. The ethynyl group is bent away from the Fe-atom and out of the plane of the adjacent carbon atoms by 2.75^{∘}. The DFT calculations and analyses of fits including ^{13}C isotopic substitution data indicate that the molecule is in an eclipsed conformation. Trends in microwave experimental values for the distance between the Fe-atom to the center of the cyclopentadienyl ring for a series of substituted ferrocenes have been analysed. This analysis provides an estimate of the gas phase distance between the Fe-atom to the center of the cyclopentadienyl ring for ferrocene of 1.65(1) Å.

aSupported by THE NATIONAL SCIENCE FOUNDATION CHE0304969